# The long way of a viscous vortex dipole

### v1

@@ A.3. Inverting the advection operator
$|\bigl{(}\nabla\mathcal{T}_{\varepsilon}\varphi\bigr{)}(\xi)|\,$ $\lesssim\,\varepsilon\int_{\{|\eta|>4\varepsilon^{-\sigma_{1}}\}}(1{+}|\xi|{+}|\eta|)\,|w(\eta)|\,{\rm d}\eta+\varepsilon^{3}\int_{\mathbb{R}^{2}}(1{+}|\xi|{+}|\eta|)^{2}|w(\eta)|\,{\rm d}\eta+|I_{\mathrm{out}}(\xi)|$
$\lesssim\,\varepsilon^{3}(1+|\xi|)^{2}\|w\|_{\mathcal{X}_{\varepsilon}}+(1+|\xi|)\exp(-c_{*}\varepsilon^{-2\sigma_{1}})\|w\|_{\mathcal{X}_{\varepsilon}}+|I_{\mathrm{out}}(\xi)|\,,$
whenever $|\xi|\leq 2\varepsilon^{-\sigma_{1}}$. Dividing both sides by $(1+|\xi|)^{3}$ and taking the $L^{q}$ norm with $q>2$, we obtain the bound involving $\nabla\mathcal{T}_{\varepsilon}\varphi$ in (4.34), in view of (A.14).
The estimate involving $\mathcal{T}_{\varepsilon}\varphi$ in (4.34) is established following exactly the same lines, and the details can thus be omitted. When $\varepsilon(|\xi|+|\eta|)\ll 1$, we now expand the kernel $K_{\varepsilon}$ itself, and not its derivative $\nabla K_{\varepsilon}$, which makes a difference because we only know that $P_{1}(\xi)=0$ under the moment assumptions in Lemma 4.6. Therefore our expansion stops at $n=1$ (instead of $n=2$ in the previous case), which explains why we only gain a factor $\varepsilon^{2}$ (instead of $\varepsilon^{3}$).
It remains to prove the estimate $\|(1+|\cdot|)\varphi\|_{L^{\infty}}\leq C\|w\|_{\mathcal{X}_{\varepsilon}}$. We already know from the proof of Lemma 4.5 that $|\varphi(\xi)|\leq C\log(2+|\xi)\|w\|_{\mathcal{X}_{\varepsilon}}$, so we can assume henceforth that $|\xi|$ is large. As $\mathrm{M}[w]=0$ by assumption, we can decompose $\varphi(\xi)=\varphi_{1}(\xi)+\varphi_{2}(\xi)$ where
$\varphi_{1}(\xi)\,=\,\frac{1}{2\pi}\int_{\{|\eta|\leq|\xi|/2\}}\log\frac{|\xi-\eta|}{|\xi|}\,w(\eta)\,{\rm d}\eta\,,\quad\varphi_{2}(\xi)\,=\,\frac{1}{2\pi}\int_{\{|\eta|>|\xi|/2\}}\log\frac{|\xi-\eta|}{|\xi|}\,w(\eta)\,{\rm d}\eta\,.$
Since $|\log(|\xi-\eta|/|\xi|)|\leq C|\eta|/|\xi|$ when $|\eta|\leq|\xi|/2$, we have $|\xi||\varphi_{1}(\xi)|\leq C\|w\|_{\mathcal{X}_{\varepsilon}}$. To bound $\varphi_{2}$, we use Hölder’s inequality and the fact that the $L^{2}$ norm of $(1+|\eta|)^{2}w(\eta)$ in the exterior domain $|\eta|>|\xi|/2$ decays rapidly as $|\xi|\to+\infty$. This gives the desired result. ∎
@@
Acknowledgements. The research of MD was supported by the Swiss State Secretariat for Education, Research and lnnovation (SERI) under contract number MB22.00034 through the project TENSE and partially supported by the GNAMPA-INdAM. The research of TG is partially supported by the grant BOURGEONS ANR-23-CE40-0014-01 of the French National Research Agency.

### v2

@@ A.4. Estimates for the stream function
$|\varphi(\xi)|\,\lesssim\,\int_{\mathbb{R}^{2}}\bigl|\log|\xi-\eta|\bigr|\,|w(\eta)|\,{\rm d}\eta\,,\qquad|\nabla\varphi(\xi)|\,\lesssim\,\int_{\mathbb{R}^{2}}\frac{1}{|\xi-\eta|}\,|w(\eta)|\,{\rm d}\eta\,.$ (A.14)
To prove the logarithmic upper bound on $|\varphi|$, we use the first inequality in (A.14) together with the crude estimates
$\bigl|\log|\xi-\eta|\bigr|\,\lesssim\,\begin{cases}|\xi-\eta|^{-1/2}&\hbox{ if }\,|\xi-\eta|\leq 1\,,\\ \log(2+|\xi|)\log(2+|\eta|)&\hbox{ if }\,|\xi-\eta|\geq 1\,.\end{cases}$
We thus find
$|\varphi(\xi)|\,\lesssim\,\int_{\{|\xi-\eta|\leq 1\}}\frac{1}{|\xi-\eta|^{1/2}}|w(\eta)|\,{\rm d}\eta\,+\,\log(2+|\xi|)\int_{\{|\xi-\eta|\geq 1\}}\log(2+|\eta|)|w(\eta)|\,{\rm d}\eta\,.$
The first integral is bounded by $C\|w\|_{L^{2}}\leq C\|w\|_{\mathcal{X}_{\varepsilon}}$, and the second one by $C\|w\|_{\mathcal{X}_{\varepsilon}}\log(2+|\xi|)$ in view of (4.25) and Hölder’s inequality. Altogether we see that $|\varphi(\xi)|\leq C\|w\|_{\mathcal{X}_{\varepsilon}}\log(2+|\xi|)$, which implies in particular that $\|(1+|\cdot|)^{-1}\varphi\|_{L^{q}}\leq C\|w\|_{\mathcal{X}_{\varepsilon}}$ for any $q>2$.
To conclude the proof of (4.31), we estimate $|\nabla\varphi|$ in (A.14) using the classical Hardy-Littlewood-Sobolev inequality, see [16, Lemma 2.1]. Given $1<p<2<q<\infty$ with $1/p=1/q+1/2$, the HLS inequality shows that $\left\lVert\nabla\varphi\right\rVert_{L^{q}}\lesssim\left\lVert w\right\rVert_{L^{p}}\lesssim\left\lVert w\right\rVert_{\mathcal{X}_{\varepsilon}}$, where in the last step we used the fact that $\mathcal{X}_{\varepsilon}\hookrightarrow L^{r}(\mathbb{R}^{2})$ for any $r\in[1,2]$.
It remains to establish (4.32). Starting from (A.14) we observe that
The last integrand is equal to $|\xi-\eta|^{-1}\bigl((1+|\eta|)|w(\eta)|^{1/2}\bigr)|w(\eta)|^{1/2}$, so we can apply the trilinear Hölder inequality with exponents $8/5$, 44, 88 to obtain
where in the last step we used the embedding $H^{1}(\mathbb{R}^{2})\hookrightarrow L^{4}(\mathbb{R}^{2})$. ∎
Proof of Lemma 4.6.
In view of Lemma 4.5, it is sufficient here to estimate $\varphi(\xi)$ and $\nabla\varphi(\xi)$ for $|\xi|\geq 1$. Using (2.6) and the assumptions on the moments of $w$, we find
$\varphi(\xi)\,=\,\frac{1}{4\pi}\int_{\mathbb{R}^{2}}\biggl\{\log\Bigl(1-\frac{2\xi\cdot\eta}{|\xi|^{2}}+\frac{|\eta|^{2}}{|\xi|^{2}}\Bigr)+\frac{2\xi\cdot\eta}{|\xi|^{2}}\biggr\}w(\eta){\rm d}\eta\,=:\,\varphi_{1}(\xi)+\varphi_{2}(\xi)\,.$
Here the decomposition $\varphi=\varphi_{1}+\varphi_{2}$ is obtained by splitting the integration domain $\mathbb{R}^{2}$ according to whether $|\eta|\leq|\xi|/4$ (for $\varphi_{1})$ or $|\eta|>|\xi|/4$ (for $\varphi_{2}$). A simple Taylor expansion shows that
$\biggl|\log\Bigl(1-\frac{2\xi\cdot\eta}{|\xi|^{2}}+\frac{|\eta|^{2}}{|\xi|^{2}}\Bigr)+\frac{2\xi\cdot\eta}{|\xi|^{2}}\biggr|\,\leq\,C\,\frac{|\eta|^{2}}{|\xi|^{2}}\,,\qquad\hbox{when}~\,|\eta|\,\leq\,\frac{|\xi|}{4}\,$,
and this readily implies that $|\xi|^{2}|\varphi_{1}(\xi)|\leq C\|w\|_{\mathcal{X}_{\varepsilon}}$. On the other hand, we observe that
$\varphi_{2}(\xi)\,=\,\frac{1}{4\pi}\int_{\{|\eta|>|\xi|/4\}}\frac{1}{|\eta|^{N}}\Bigl(\log|\xi-\eta|^{2}-\log|\xi|^{2}+\frac{2\xi\cdot\eta}{|\xi|^{2}}\Bigr)|\eta|^{N}w(\eta){\rm d}\eta\,$,
where the integer $N\in\mathbb{N}$ is arbitrary. Proceeding as in the proof of Lemma 4.5 and using the fact that $|\eta|^{N}w(\eta)$ decays rapidy as $|\eta|\to\infty$, we deduce that $|\xi|^{N}|\varphi_{2}(\xi)|\leq C\|w\|_{\mathcal{X}_{\varepsilon}}\log(2+|\xi|)$. Taking $N>2$, we conclude that $(1+|\xi|)^{2}|\varphi(\xi)|\leq C\|w\|_{\mathcal{X}_{\varepsilon}}$ as asserted in (4.33).
We proceed similarly to estimate $\nabla\varphi(\xi)$ for $|\xi|\geq 1$. Our starting point is the formula
$\nabla\varphi(\xi)\,=\,\frac{1}{2\pi}\int_{\mathbb{R}^{2}}\biggl\{(\xi-\eta)\Bigl(\frac{1}{|\xi-\eta|^{2}}-\frac{1}{|\xi|^{2}}\Bigr)-\frac{2\xi\cdot\eta}{|\xi|^{4}}\,\xi\biggr\}w(\eta){\rm d}\eta\,=:\,\psi_{1}(\xi)+\psi_{2}(\xi)\,$,
where $\psi_{1}(\xi),\psi_{2}(\xi)$ correspond to integrating over the regions $|\eta|\leq|\xi|/4$ and $|\eta|>|\xi|/4$, respectively. Again, a Taylor expansion shows that
$\biggl|(\xi-\eta)\Bigl(\frac{1}{|\xi-\eta|^{2}}-\frac{1}{|\xi|^{2}}\Bigr)-\frac{2\xi\cdot\eta}{|\xi|^{4}}\,\xi\biggr|\,\leq\,C\,\frac{|\eta|^{2}}{|\xi|^{3}}\,,\qquad\hbox{when}~\,|\eta|\,\leq\,\frac{|\xi|}{4}\,$,
so that $|\xi|^{3}|\psi_{1}(\xi)|\leq C\|w\|_{\mathcal{X}_{\varepsilon}}$. On the other hand, given any $N\in\mathbb{N}$, we have
$|\psi_{2}(\xi)|\,\leq\,C\int_{\{|\eta|>|\xi|/4\}}\frac{1}{|\eta|^{N}}\Bigl(\frac{1}{|\xi-\eta|}+\frac{|\eta|}{|\xi|^{2}}\Bigr)|\eta|^{N}w(\eta){\rm d}\eta\,.$
Proceeding as in the proof of Lemma 4.5, we deduce that
$\|\,|\cdot|^{N}\psi_{2}\|_{L^{q}}\,\leq\,C\|w\|_{\mathcal{X}_{\varepsilon}}\,,\quad\hbox{and}\quad\|\,|\cdot|^{N}\psi_{2}\|_{L^{\infty}}\,\leq\,C\bigl(\|\nabla w\|_{\mathcal{X}_{\varepsilon}}^{1/2}+\|w\|_{\mathcal{X}_{\varepsilon}}^{1/2}\bigr)\|w\|_{\mathcal{X}_{\varepsilon}}^{1/2}\,.$
When $N=2$, the first estimate also holds for $\psi_{1}$, hence for $\nabla\varphi=\psi_{1}+\psi_{2}$, which concludes the proof of (4.33). Similarly, the second estimate also holds for $\psi_{1}$ when $N=3$, which implies that $\nabla\varphi$ satisfies (4.34). ∎
Proof of Corollary 4.7.
Denoting $\eta=\widetilde{\xi}-\varepsilon^{-1}e_{1}$, we have $\bigl(\mathcal{T}_{\varepsilon}\varphi\bigr)(\xi)=\varphi(\eta)$ by (2.10). We assume here that $\xi\in\mathrm{I}_{\varepsilon}$, so that $|\xi|\leq 2\varepsilon^{-\sigma_{1}}\ll\varepsilon^{-1}$ if $\varepsilon$ is sufficiently small. This in turn implies that $(2\varepsilon)^{-1}\leq|\eta|\leq 2\varepsilon^{-1}$. In particular, we have
$\bigl|\mathbbm{1}_{\mathrm{I}_{\varepsilon}}(\xi)\bigl(\mathcal{T}_{\varepsilon}\varphi)(\xi)\bigr|\,=\,\Bigl|\mathbbm{1}_{\mathrm{I}_{\varepsilon}}(\xi)\frac{1}{(1+|\eta|)^{2}}\,(1+|\eta|)^{2}\varphi(\eta)\Bigr|\,\leq\,C\varepsilon^{2}\bigl\|(1+|\cdot|)^{2}\varphi\bigr\|_{L^{\infty}}\,\leq\,C\varepsilon^{2}\|w\|_{\mathcal{X}_{\varepsilon}}\,$,
where the last step follows from (4.33). This proves the first inequality in (4.35). On the other hand, the proof of Lemma 4.6 shows that $\nabla\varphi=\psi_{1}+\psi_{2}$ where
$\bigl\|(1+|\cdot|)^{3}\psi_{1}\bigr\|_{L^{\infty}}+\bigl\|(1+|\cdot|)^{3}\psi_{2}\bigr\|_{L^{q}}\,\leq\,C\|w\|_{\mathcal{X}_{\varepsilon}}\,$,
for any $q>2$. As a consequence, we have
$\bigl\|\mathbbm{1}_{\mathrm{I}_{\varepsilon}}(\xi)(1+|\xi|)^{-1}\psi_{1}(\eta)\bigl\|_{L^{q}}\,$ $\leq\,\bigl\|\mathbbm{1}_{\mathrm{I}_{\varepsilon}}(\xi)\frac{1}{(1+|\eta|)^{3}}\,(1+|\eta|)^{3}\psi_{1}(\eta)\bigl\|_{L^{\infty}}\,\leq\,C\varepsilon^{3}\|w\|_{\mathcal{X}_{\varepsilon}}\,$,
$\bigl\|\mathbbm{1}_{\mathrm{I}_{\varepsilon}}(\xi)(1+|\xi|)^{-1}\psi_{2}(\eta)\bigl\|_{L^{q}}\,$ $\leq\,\bigl\|\mathbbm{1}_{\mathrm{I}_{\varepsilon}}(\xi)\frac{1}{(1+|\eta|)^{3}}\,(1+|\eta|)^{3}\psi_{2}(\eta)\bigl\|_{L^{q}}\,\leq\,C\varepsilon^{3}\|w\|_{\mathcal{X}_{\varepsilon}}\,.$
Since $\nabla_{\xi}\bigl(\mathcal{T}_{\varepsilon}\varphi\bigr)(\xi)=\widetilde{\nabla}_{\eta}\varphi(\eta)=\widetilde{\psi_{1}}(\eta)+\widetilde{\psi_{2}}(\eta)$, this implies the second inequality in (4.35). ∎
@@
Acknowledgements. The research of MD was supported by the Swiss State Secretariat for Education, Research and lnnovation (SERI) under contract number MB22.00034 through the project TENSE and by the Swiss National Science Foundation (SNF Ambizione grant PZ00P2_223294). The research of TG is partially supported by the grant BOURGEONS ANR-23-CE40-0014-01 of the French National Research Agency.